OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that contains the following points? (-2,-20),(0,-4),(4,-20)

OpenStudy (anonymous):

y = -2x^2 + 4x - 4

OpenStudy (anonymous):

How did you get that @-EtherealEbullience ?

OpenStudy (anonymous):

y = ax^2 + bx + c, -20 = 4a - 2b + c -4 = c -20 = 16a + 4b + c

OpenStudy (anonymous):

For the third equation, how did you get 16a?

OpenStudy (anonymous):

4x^2

OpenStudy (anonymous):

!!! Thank you ! (: